OpenStudy (anonymous):

Arnold's company reimburses his expenses on food, lodging, and conveyance during business trips. The company pays $40 a day for food and lodging and $0.50 for each mile traveled. Arnold drove 400 miles and was reimbursed $2600. Part A: Create an equation that will determine the number of days x on the trip Part B: Solve this equation justifying each step with an algebraic property of equality. Part C: How many days did Arnold spend on this trip?

OpenStudy (anonymous):

40d + .50(400) = 2600 -- your equation 40d + 200 = 2600 40d = 2600 - 200 (subtraction equality) 40d = 2400 d = 2400/40 (division equality) d = 60 -- He spent 60 days any questions ?

OpenStudy (anonymous):

Thank you so much !

OpenStudy (anonymous):

no problem :)